OpenStudy (anonymous):

solve the equation (x-8)(2x+7=0

OpenStudy (agreene):

x= -7/2 and x=8 start by FOILing and then do quadratic or whatever method you prefer.

jhonyy9 (jhonyy9):

x-8=0 x=8 2x+7=0 2x=-7 x= -7/2